OpenStudy (anonymous):

how do you simplify 5(x+4)+6(x+3)

OpenStudy (aroub):

do you know how to distribute?

OpenStudy (anonymous):

no not really.... im new to all of this....

OpenStudy (anonymous):

5x+5*4 + 6x + 6*3 -> 11x + 38

OpenStudy (anonymous):

5x+20+6x+18=11x+38

OpenStudy (anonymous):

how do you distribute???

OpenStudy (anonymous):

for example 5*(2+3) you multiply every term inside the bars : 5*2 + 5*3

OpenStudy (anonymous):

thanks i think i understand a bit bettr now. im still gonna need to keep up on studies for it

OpenStudy (aroub):

okay, you have 5(x+4)+6(x+3) all you have to do is just multiply the 5 by the numbers inside the brackets like this: 5(x)+ 5(4)= 5x+20 and same thing for 6 you multiply it by everything inside the brackets 6(x) +6(3)=6x+18 now you add the numbers you got 5x+20+6x+18 11x+ 38

OpenStudy (anonymous):

thank you soooooooo much.... so i need help on another one first i wanna see if i get right...

OpenStudy (aroub):

okay :)

OpenStudy (anonymous):

problem is 2(x-4y+1) + 3(-2x+y-2) imma try and figure out the answer first

OpenStudy (aroub):

ok and tell me what will you get so i could check for you.. :)

OpenStudy (anonymous):

i think i got the first part right the second part was a lil confusing.... for the first part i got 2x-8y+2

OpenStudy (aroub):

correct! :) , do the same thing for the second part

OpenStudy (anonymous):

i did but it came up as a multipplication

OpenStudy (aroub):

ok, just tell me what did you get .. please

OpenStudy (anonymous):

-21 3*-6x+3y-6

OpenStudy (aroub):

3(-2x+y-2) 3(-2x)+3(y)+3(-2) -6x+3y-6 ^ that's the true answer :)

OpenStudy (aroub):

now add them..